OpenStudy (anonymous):

Just another cute problem, How many eight digit numbers formed using the digit \( \{7,8,9\} \) has exactly three \( 7 \)'s ?

OpenStudy (mr.math):

If I get the question right then the answer should be \(8\cdot 7\cdot 6\cdot {5\choose 2}\).